# Artificial Intelligence Exploring the Patent Field

### v1

@@ 4.1 Data Sources
Patent applications are submitted to and granted by patent offices. To stimulate innovation and serve the society, patent offices provide detailed information about existing patents, patent applications, and the legal status of patents—previously on paper, nowadays online. Large patent offices include the United States Patent and Trademark Office (USPTO) and the European Patent Office (EPO). These offices also provide access to download large-scale bulk datasets or tools to explore and analyze patent data. For example, PatentsView is a platform developed by the USPTO, providing accessible and user-friendly interfaces to explore US patent data, with various tools for visualization and analysis. Apart from patent offices, there are searchable databases that contain patents from multiple countries and offices, such as Google Patents. We list a broad range of patent offices and databases in Table 1.
@@
Patent Office & Database Brief Description
USPTO
United States Patent and Trademark Office is a comprehensive source for US patents and applications. https://www.uspto.gov/
EPO
European Patent Office offers a variety of tools, such as espacenet, for searching European and also foreign patents, which the tools can link to patent families. https://www.epo.org/en
WIPO
World Intellectual Property Organization includes global patent documents and applications from participating national and regional patent offices. https://www.wipo.int/portal/en/index.html
CNIPA / SIPO
State Intellectual Property Office, formerly China National Intellectual Property Administration, provides a database for searching Chinese patents. http://english.cnipa.gov.cn/
JPO
Japan Patent Office provides information on Japanese patents and utility models. https://www.jpo.go.jp/e/
UK IPO
United Kingdom Intellectual Property Office gives access to UK, Jersey, and Guernsey patents, . https://www.ipo.gov.uk/
KIPO
Korea Intellectual Property Office includes the Korean Patent System and Utility Model System. https://www.kipo.go.kr/
Auspat
The Australian patent database. https://www.ipaustralia.gov.au/
Google Patents
This is a free and searchable database of patents and patent applications from multiple countries. https://patents.google.com/
PatentsView
A platform developed by the USPTO, providing interfaces for exploring U.S. patent data. https://patentsview.org/
Espacenet
Espacenet allows free access to over 140 million patent documents provided by EPO. https://worldwide.espacenet.com/
Patentscope
Patentscope belongs to WIPO, allowing searches in international and national patent collections. https://www.wipo.int/patentscope/en/
Global Dossier
Global Dossier is developed and maintained by a collaboration of USPTO, EPO, JPO, KIPO, and SIPO to provide a shared database and access point to IP rgisters and th corresponding procedural files. https://globaldossier.uspto.gov
Lens
Lens integrates scholarly and patent knowledge, providing patent search and analysis tools. https://www.lens.org/
@@ 4.2 Curated Data Collections
Datasets. Patent offices provide large-scale raw data in the patent domain. Developers and researchers rely on well-curated datasets for informing development work and tools exploiting the contained knowledge. We summarize representative curated patent datasets in Table 2. We aim to reduce the time and effort for data searching in the patent domain by presenting these ready-to-use datasets, especially the publicly available ones.
Table 2: Curated patent datasets.
@@
USPTO-2M [17] USPTO Classification 2M
Utility patent documents classified into 637 classes. Original link: http://mleg.cse.sc.edu/DeepPatent/
USPTO-3M [18] USPTO Classification 3M
Collected from Google Patents Public Datasets using SQL. Data information at: https://doi.org/10.1016/j.wpi.2020.101965
WIPO-Alpha [19] WIPO Classification 75K
Documents classified into 114 classes and 451 subclasses. Original link: https://www.wipo.int/ibis/datasets
@@
Industry dataset [20]
Classification 10K
A patent dataset focusing on the crop industry domain. Available at: https://github.com/GT4SD/domain-adaptive-patent-classifier
DeepPatent [21] USPTO
@@
45K
Image-based patent retrieval benchmark. Available at: https://github.com/GoFigure-LANL/DeepPatent-dataset
PatentMatch [22] EPO
@@
6.3M
Pairs of claims and text passages from cited patents. Available at: https://hpi.de/naumann/s/patentmatch
[23]
@@
2.5M
Patent pairs with labels of cited/random. Available at: https://github.com/helmersl/patent_similarity_search
TFH [24] USPTO
@@
40K
Labeled patent dataset for entities and the semantic relations extraction. Available at: https://github.com/awesome-patent-mining/TFH_Annotated_Dataset
[25]
@@
5.5M
Granted and rejected patent applications. Available at: https://github.com/Catchy1997/DLPAP
[27] USPTO
@@
117K
Triples of a plaintiff, defendant, and a patent. Available at: https://github.com/USTCwuhan/MANTF-master
BigPatent [28] USPTO Summarization 1.3M
Patent documents with human-written abstractive summaries. Available at: https://evasharma.github.io/bigpatent/
[29] Simplification 0.2M
Silver standard dataset for patent simplification. Available at: https://github.com/slvcsl/patentSilverStandard
GED [30] EPO Translation 19K
Patent language pairs French–English, and German–English. Available at: https://huggingface.co/datasets/mwirth-epo/epo-nmt-datasets
EuroPat [31] EPO Translation 51M
6 official European languages paired with English. Available at: https://opus.nlpl.eu/EuroPat.php
[32] USPTO Patent writing 0.6M
Patent claims of the granted U.S. utility patents in 2013. Data information at: https://doi.org/10.1016/j.wpi.2020.101983
HUPD [9] USPTO Multi-purpose 4.5M
A large-scale patent corpus containing 34 data fields. Available at: https://patentdataset.org/
The number of curated datasets for patent classification and patent retrieval is typically larger than others because the data collection process is simple. Each granted patent is assigned classification codes and contains referenced patents. Hence, researchers can formulate the datasets by collecting and filtering patents without further labeling. For patent novelty quantification and prediction, Arts et al. considered patents connected to awards such as the Nobel Prize as novel because they radically impacted technological progress and patenting [25]. In contrast, patents were considered lacking novelty if they were granted by the United States Patent and Trademark Office but simultaneously rejected by both the European Patent Office and Japan Patent Office. However, this collection deviates from the formal definition of novelty introduced in Section 7.4.1. For patent simplification, there is only a silver standard dataset [29]. In text generation tasks, texts written by humans are usually considered the gold standard. Since patent simplification requires extensive expertise and effort, it is expensive and time-consuming to obtain a gold standard for patent simplification. Thus, Casola et al. adopted automated tools to generate simpler texts for patents and named it the silver standard [29].
Notably, the Harvard USPTO Patent Dataset (HUPD) [9] is a recently presented large-scale multi-purpose dataset. It contains more than 4.5 million patent documents with 34 data fields, providing opportunities for various tasks. The paper demonstrated four types of usages of this dataset, including granting prediction, subject classification, language modeling, and summarization.
@@
The intellectual property arm of the Conference and Labs of the Evaluation Forum (CLEF-IP) [^2] CLEF: https://www.clef-initiative.eu/ focuses on evaluation tasks related to intellectual property, particularly in patent retrieval and analysis. Each task usually contains a curated patent dataset for desired aims, including various information such as text, images, and metadata [33].
@@
The Japanese National Institute of Informatics Testbeds and Community for Information access Research (NTCIR) [^3] NTCIR: https://research.nii.ac.jp/ntcir/index-en.html provides datasets and organizes shared tasks to facilitate research in information retrieval, natural language processing, and related areas. Patent-related tasks at the NTCIR range from patent retrieval and classification to text mining and machine translation [34].

### v2

@@ 4.1 Data Sources
Patent applications are submitted to and granted by patent offices. To stimulate innovation and serve the society, patent offices provide detailed information about existing patents, patent applications, and the legal status of patents—previously on paper, nowadays online. Large patent offices include the United States Patent and Trademark Office (USPTO) and the European Patent Office (EPO). These offices also provide access to download bulk datasets or tools to explore and analyze patent data. For example, PatentsView is a platform developed by the USPTO, providing accessible and user-friendly interfaces to explore US patent data, with various tools for visualization and analysis. Apart from patent offices, there are searchable databases that contain patents from multiple countries and offices, such as Google Patents. We list a broad range of patent offices and databases in Table 1.
Patent Office & Database Brief Description
USPTO United States Patent and Trademark Office is a comprehensive source for US patents and applications. https://www.uspto.gov/
EPO European Patent Office offers a variety of tools, such as Espacenet, for searching European and also foreign patents, which the tools can link to patent families. https://www.epo.org/en
WIPO World Intellectual Property Organization includes global patent documents and applications from participating national and regional patent offices. https://www.wipo.int/portal/en/index.html
CNIPA / SIPO State Intellectual Property Office, formerly China National Intellectual Property Administration, provides a database for searching Chinese patents. http://english.cnipa.gov.cn/
JPO Japan Patent Office provides information on Japanese patents and utility models. https://www.jpo.go.jp/e/
UK IPO United Kingdom Intellectual Property Office gives access to UK, Jersey, and Guernsey patents, . https://www.ipo.gov.uk/
KIPO Korea Intellectual Property Office includes the Korean Patent System and Utility Model System. https://www.kipo.go.kr/
Auspat The Australian patent database. https://www.ipaustralia.gov.au/
Google Patents This is a free and searchable database of patents and patent applications from multiple countries. https://patents.google.com/
PatentsView A platform developed by the USPTO, providing interfaces for exploring U.S. patent data. https://patentsview.org/
Espacenet Espacenet allows free access to over 140 million patent documents provided by EPO. https://worldwide.espacenet.com/
Patentscope Patentscope belongs to WIPO, allowing searches in international patent collections. https://www.wipo.int/patentscope/en/
Global Dossier Global Dossier is developed and maintained by a collaboration of USPTO, EPO, JPO, KIPO, and SIPO to provide a shared database and access point to IP registers and the corresponding procedural files. https://globaldossier.uspto.gov
Lens Lens integrates scholarly and patent knowledge, providing patent search and analysis tools. https://www.lens.org/
@@ 4.2 Curated Data Collections
USPTO-2M (Li et al., 2018) USPTO Classification 2M Utility patent documents classified into 637 classes. Original link: http://mleg.cse.sc.edu/DeepPatent/
USPTO-3M (Lee and Hsiang, 2020b) USPTO Classification 3M Collected from Google Patents Public Datasets using SQL. Data information at: https://doi.org/10.1016/j.wpi.2020.101965
WIPO-Alpha (Fall et al., 2003) WIPO Classification 75K Documents classified into 114 classes and 451 subclasses. Original link: https://www.wipo.int/ibis/datasets
@@
Industry dataset (Christofidellis et al., 2023)
Classification 10K A patent dataset focusing on the crop industry domain. Available at: https://github.com/GT4SD/domain-adaptive-patent-classifier
DeepPatent (Kucer et al., 2022) USPTO
@@
45K Image-based patent retrieval benchmark. Available at: https://github.com/GoFigure-LANL/DeepPatent-dataset
PatentMatch (Risch et al., 2020a) EPO
@@
6.3M Pairs of claims and text passages from cited patents. Available at: https://hpi.de/naumann/s/patentmatch
Helmers et al. (2019)
@@
2.5M Patent pairs with labels of cited/random. Available at: https://github.com/helmersl/patent_similarity_search
TFH (Chen et al., 2020) USPTO
@@
40K Labeled patent dataset for entities and the semantic relations extraction. Available at: https://github.com/awesome-patent-mining/TFH_Annotated_Dataset
Arts et al. (2021)
@@
5.5M Granted and rejected patent applications. Available at: https://github.com/Catchy1997/DLPAP
Wu et al. (2023) USPTO
@@
117K Triples of a plaintiff, defendant, and a patent. Available at: https://github.com/USTCwuhan/MANTF-master
BigPatent (Sharma et al., 2019) USPTO Summarization 1.3M Patent documents with human-written abstractive summaries. Available at: https://evasharma.github.io/bigpatent/
Casola et al. (2023) Simplification 0.2M Silver standard dataset for patent simplification. Available at: https://github.com/slvcsl/patentSilverStandard
GED (Wirth et al., 2023) EPO Translation 19K Patent language pairs French–English, and German–English. Available at: https://huggingface.co/datasets/mwirth-epo/epo-nmt-datasets
EuroPat (Heafield et al., 2022) EPO Translation 51M 6 official European languages paired with English. Available at: https://opus.nlpl.eu/EuroPat.php
Lee and Hsiang (2020a) USPTO Patent writing 0.6M Patent claims of the granted U.S. utility patents in 2013. Data information at: https://doi.org/10.1016/j.wpi.2020.101983
HUPD (Suzgun et al., 2023) USPTO Multi-purpose 4.5M A large-scale patent corpus containing 34 data fields. Available at: https://patentdataset.org/
Table 2: Curated patent datasets.
Datasets. Patent offices provide large-scale raw data in the patent domain. Developers and researchers rely on well-curated datasets for development and research. We summarize representative curated patent datasets in Table 2. We aim to reduce the time and effort for data searching in the patent domain by presenting these ready-to-use datasets, especially the publicly available ones.
The number of curated datasets for patent classification and patent retrieval is typically larger than others because the data collection process is simple. Each granted patent is assigned classification codes and contains referenced patents. Hence, researchers can formulate the datasets by collecting and filtering patents without further labeling. For patent novelty quantification and prediction, Arts et al. (2021) considered patents connected to awards such as the Nobel Prize as novel because they radically impacted technological progress and patenting. In contrast, patents were considered lacking novelty if they were granted by the United States Patent and Trademark Office but simultaneously rejected by both the European Patent Office and Japan Patent Office. However, this collection may deviate from the formal definition of novelty introduced in Section 5.4.1. For patent simplification, there is only a silver standard dataset (Casola et al., 2023). In text generation tasks, texts written by humans are usually considered the gold standard. Since patent simplification requires extensive expertise and effort, it is expensive and time-consuming to obtain a gold standard for patent simplification. Thus, Casola et al. (2023) adopted automated tools to generate simpler texts for patents and named it the silver standard.
Notably, the Harvard USPTO Patent Dataset (HUPD) (Suzgun et al., 2023) is a recently presented large-scale multi-purpose dataset. It contains more than 4.5 million patent documents with 34 data fields, providing opportunities for various tasks. The corresponding paper demonstrates four types of usages of this dataset, including granting prediction, subject classification, language modeling, and summarization.
@@
The intellectual property arm of the Conference and Labs of the Evaluation Forum (CLEF-IP) [^9] CLEF: https://www.clef-initiative.eu/ focuses on evaluation tasks related to intellectual property, particularly in patent retrieval and analysis. Each task usually contains a curated patent dataset for desired aims, including various information such as text, images, and metadata (Piroi and Hanbury, 2017).
@@
The Japanese National Institute of Informatics Testbeds and Community for Information access Research (NTCIR) [^10] NTCIR: https://research.nii.ac.jp/ntcir/index-en.html provides datasets and organizes shared tasks to facilitate research in information retrieval, natural language processing, and related areas. Patent-related tasks at the NTCIR range from patent retrieval and classification to text mining and machine translation (Lupu et al., 2017).